You catch an exotic beetle while you are in the garden. You tie one of the feet of the beetle to string and then hold the other end of string. Once you release the beetle from your hand, the beetle starts flying in circle in a counterclockwise manner with a velocity of 8 m/s2. The length of the string from your hand to the beetle’s foot is 14 inches. What is the angular momentum of the beetle?
1
Expert's answer
2020-04-05T14:32:50-0400

m is the mass of the beetle in kg.

 The angular momentum of the beetle:


"M=mvr=(8)(14\cdot 0.3048)m=(34.14m)\frac{kg\cdot m^2}{s}"

Direction: counterclockwise.